OpenStudy (anonymous):

A trapezoid has base lengths of (6x - 1) units and 3 units. Its midsegment has a length of (5x - 3) unit. What is the value of x? What does it mean by units?

OpenStudy (janu16):

so you dont need help solvong it ?

OpenStudy (anonymous):

Hmm, well, I don't know what to do with that information.

OpenStudy (janu16):

you said you want to now what units is

OpenStudy (janu16):

cant help you trying to explan tht cause i don tknow how to explain but i can help you solve it

OpenStudy (anonymous):

No that's the question, I have no idea what it means by units and I drew that as well but I don't know if to go with it this way: 6x - 1 = 5x - 3 OR I don't know.

OpenStudy (anonymous):

But if you can help me solve it, I can sort of understand better (works for me mot of hte time)

OpenStudy (janu16):

it goes 5x-3=1/2(6x-1+3)

OpenStudy (anonymous):

Okay because them midsegment, that's why you have 1/2 in there right

OpenStudy (janu16):

yes so now solve this (6x-1+3)

OpenStudy (janu16):

what -1+3

OpenStudy (anonymous):

I got x = 2

OpenStudy (janu16):

now we have 5x-3=1/2(6x+2)

OpenStudy (janu16):

divide (6x+2) with 1/2

OpenStudy (anonymous):

Like I said, I got x = 2, is that right?

OpenStudy (janu16):

yes i tohught you were talking about something else but yes

OpenStudy (janu16):

@moonrose

OpenStudy (anonymous):

Oh okay! Thank you so much for your help, the 1/2 didn't cross my mind and I just have a hard time locating where to put the info given to me in the diagram.

OpenStudy (janu16):

yes np